in HF the experimental bond length value is less than the sum of covalent radii?
1
Expert's answer
2015-09-08T03:23:16-0400
The experimental bond length for HF was found to be 0.917 angstroms on the NIST website. https://www.uwosh.edu/faculty_staff/gutow/Chem_371_S09/SNRE%20Webpage/HF%20Calculations/HF%20Calculations.html
The sum of covalent radii 73pm. R=73*10^-12 m = 7.2*10^-11 m.
So, in HF the experimental bond length value is less than the sum of covalent radii.